OpenStudy (anonymous):

I need help! Is the relationship shown by the data linear? If so, model the data with an equation. x | y -7 5 -5 9 -3 13 -1 17

OpenStudy (solomonzelman):

try this -7/5 = -7/5 and -5 /9 = -5/9 is =-5/9=-7/5 ? if that's not true the variation is indirect.

OpenStudy (solomonzelman):

You can tell it is not a direct variation right?

OpenStudy (anonymous):

right..

OpenStudy (solomonzelman):

Good!

OpenStudy (anonymous):

so is it the relationship is not linear.

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

What do you think?

OpenStudy (anonymous):

i think it might be that. but i dont want to get it wrong..

OpenStudy (solomonzelman):

Well, you can see that it's not "linear" if you know what that means, and you do...

OpenStudy (anonymous):

hmmm..

OpenStudy (callisto):

@SolomonZelman May I ask if linear relationship the same as direct variation?

OpenStudy (callisto):

To my understanding, direct variation (y ∝ x) means that y = kx, where k is a constant. However, for linear relationship, it can be represented in a linear equation, that is y=mx+c, where m and c are constants. In this case, even if their ratios are not the same, it does not imply that they are not in a linear relationship. If the definition I have is correct, one of the way to solve this problem may be as follows: 1) Establish the equation y = mx + c 2) Plug two sets of data into the equation to solve m and c, which are constants 3) Plug x in the 3rd and 4th pair of data into mx+c (in this step, you should know the value of m and c), and see if it equals to the value of their y's. If they equals to their y's, then they have a linearly relationship, and vice versa.

OpenStudy (anonymous):

was he wrong?

OpenStudy (callisto):

Have you solved it?